OpenStudy (anonymous):

The length of one leg of an isosceles right triangle is 3 ft. What is the perimeter of the triangle?

OpenStudy (anonymous):

Pythagorean theorem can be used when finding a missing side of a right triangle. So if the lengths of the sides of your triangle fit the Pythagorean theorem, then it should be 3, 4, and 5. So 3+4+5=12

OpenStudy (anonymous):

WHAT you say is the correct answer of all these? \[3+3\sqrt{2}\]

OpenStudy (anonymous):

\[6+3\sqrt{2}\]

OpenStudy (anonymous):

\[3+3\sqrt{3}\]

OpenStudy (anonymous):

or

OpenStudy (anonymous):

\[6+3\sqrt{3}\]

OpenStudy (anonymous):

6+ 3√2

OpenStudy (anonymous):

In the right triangle below, tanA = 0.45. What is the approximate length of AB?

OpenStudy (anonymous):

I'm sorry but I'm not sure how to do that

OpenStudy (anonymous):

In a right triangle, angle A measures 20°. The side opposite angle A is 10 centimeters long. Approximately how long is the hypotenuse of the triangle?

OpenStudy (anonymous):

29.2

OpenStudy (anonymous):

What is the approximate length of arc s on the circle below? Use 3.14 for

OpenStudy (anonymous):

What do you need to find?

OpenStudy (anonymous):

What is the approximate length of arc s on the circle below? Use 3.14 for mc024-1.jpg. Round your answer to the nearest tenth.

OpenStudy (anonymous):

Do you have choices?